OpenStudy (anonymous):

Solve for: 4x^2 - 9 = 0

OpenStudy (anonymous):

Add 9 to both sides to get: 4x^2 = 9 Divide both sides by 4 to get: x^2 = 9/4. Remember that a square root can take both values, a negative and a positive, so we are left with: x = +- sqrt(9/4) = +- (3/2)